show that the least value of the sum of a positive number and its reciprocal is 2.
 
 
please answer the above problem as soon as possible.

Grade:11

1 Answers

Susmita
5 years ago
Let a be the number.Its reciprocal is 1/a.
Sum s=a+(1/a)
For least possible value
ds/da=0
Or,1-(1/a2)=0
Or,a=\pm 1
We shall discard a=-1 as the question say positive number.
a=1
S=1+1=2
If it helps please approve.thank u.